# Optimal quantum transport on a ring via locally monitored chiral quantum walks

### v1

@@ IV.2 Dark states
$\hat{U}(\tau)|\gamma_{n,m}\rangle=$ $\sqrt{\frac{N}{2}}e^{-i\lambda_{m}\tau}\Big{(}\langle\delta|\lambda_{n}\rangle|\lambda_{m}\rangle$
$-\langle\delta|\lambda_{m}\rangle e^{i(\lambda_{m}-\lambda_{n})\tau}|\lambda_{n}\rangle\Big{)},$ (11)
@@
in which case the system never visits |$\ket{\delta}$, resulting in $P_{\rm det}(\infty)=0$.
If no pairs of eigenvalues $\lambda_{m},\lambda_{n}$ satisfy the condition (12) and $\langle\lambda_{j}|\delta\rangle\neq 0$ $\forall j$, then the Hilbert space is fully bright, and there are no dark states [38]. Note that when Eq. (12) holds true, we have that $[\mathbb{I}-\hat{D}]\hat{U}(\tau)|\gamma_{n,m}\rangle\,=\,e^{-i\lambda_{m}\tau}|\gamma_{n,m}\rangle$, i.e. the Perron–Frobenius operator $\hat{O}$ acts on dark states solely by multiplying them with a phase factor; that is, dark states are eigenstates of $\hat{O}$ with eigenvalues of unit modulus. Moreover, they are the only possible eigenstates of modulus 1 of $\hat{O}$. [^4] Indeed, since $\hat{U}(\tau)$ is a unitary operator, it preserves the norm. Therefore, for Perron-Frobenius operator to have an eigenvalue of modulus 1, it must be that $\hat{U}(\tau)|\mu\rangle$ (with $|\mu\rangle$ eigenstate corresponding to a unit modulus eigenvalue of $\hat{O}$) is orthogonal to $|\delta\rangle$. Otherwise, the action of $[\mathbb{I}-\hat{D}]$ would reduce its norm, making it impossible for the corresponding eigenvalue to have modulus 1. The only way $\hat{O}$ can act on an eigenstate with eigenvalue of modulus 1 is via the time evolution operator followed by the identity. Consequently, such an eigenstate must be an eigenstate of the time evolution operator $\hat{U}(\tau)$ (which acts by multiplication with a phase) and simultaneously unaffected by $[\mathbb{I}-\hat{D}]$, meaning that the latter acts as the identity on it.
In our model the condition $\langle\lambda_{j}|\delta\rangle\neq 0\;\forall j$ is satisfied by any localized state $|\delta\rangle$ [see Eq. (2)], as our detection states, while the condition (12) can be satisfied in two ways:
@@
For even $N$, when $\phi=0$ there exist $\frac{N}{2}-1$ two-fold degenerate levels. In this case, all the stationary dark states constructed from the pairs of eigenstates in each degenerate level are orthogonal to the initial state. Therefore, they neither contribute to the dynamics of the system nor affect the detection probability. When $\phi=\pm\pi/N$ there exist $N/2$ two-fold degenerate levels. It can be proved (see Appendix B) that the initial state lies in the dark subspace spanned by the stationary dark states constructed from these degenerate levels. As a consequence, in this case the walker is never detected in $|\delta\rangle$, yielding to $P_{\rm det}(\pm\pi/N)=0$ in Fig. 2(a).
For odd $N$, calculations analogous to those in Appendix B reveal that for $\phi=0,\pm\pi/N$ the total overlap of the initial state with the dark space is $1/2$, thus explaining the minima of the detection probability for these values of $\phi$ in Fig. 2(b). For such values of $\phi$, there exist $(N-1)/2$ two-fold degenerate levels, each contributing a dark state that, together, form the basis spanning the dark space.
(ii) Using any nondegenerate couple of eigenvalues $\lambda_{m}\neq\lambda_{n}$, there may exist proper combinations of $\tau$ and $\phi$ that lead to $\lambda_{m}\tau=\lambda_{n}\tau+2k\pi$ for some $k\in\mathbb{Z}$. Plugging the eigenvalues (3) in the latter expression, i.e. Eq. (12), we derive the condition relating $\tau$ and $\phi$ under which dark states appear
@@
Such a lower bound is exactly attained for even $N$ when $\phi=0$. In all other cases, the $\tau$ at which the first dark state appears approaches $\pi/2$ as $N$ grows. This is because the discrete arguments of the sine functions in Eq. (13) become denser in the interval $[0,2\pi)$ for increasing $N$. For large $N$, it is therefore possible to find pairs $(m,n)$ such that the product of the two sine functions, regardless of $\phi$, is arbitrarily close to 1, implying the condition $\tau\gtrsim\pi/2$. In other words, for any $\phi$ the first dark state associated with nondegenerate levels will arise at $\tau\to\pi/2$ as $N\to\infty$, appearing as a dark-state line at $\tau\to\pi/2$ in the parameter space $(\phi,\tau)$. In general, for finite $N$, solutions to Eq. (13) constitute dark-state curves in the parameter space $(\phi,\tau)$, which correspond to low-detection-probability curves. Figures 2(a,b) represent a clear example: The detection probability in the parameter space $(\phi,\tau)$ is low along the curves associated with the presence of dark states, including a nearly straight line at $\tau\gtrsim\pi/2$. In the region $\tau\geq\pi/2$, the dark-state curves become denser as $N$ increases, because the number of dark states from nondegenerate levels is of order $O(N^{2})$, and also as $\tau$ increases (see Appendix C).
@@
To summarize, in the region $\tau<\pi/2$, for even $N$ dark states at $\phi=0$ do not affect the evolution of $|\psi_{0}\rangle$ yielding $P_{\rm det}(\infty)=1$, while for $\phi=\pm\pi/N$ the initial state is a dark state, yielding $P_{\rm det}(\infty)=0$; for odd $N$, the initial state has a finite overlap with dark states at $\phi=0,\pm\pi/N$, yielding $P_{\rm det}(\infty)=1/2$. For $\phi\neq 0,\pm\pi/N$, the Hilbert space is fully bright in this region, so $P_{\rm det}(\infty)=1$ for both even and odd $N$. On the other hand, in the region $\tau\geq\pi/2$, the number of dark states increases as $N$ and $\tau$ increase, yielding low-detection-probability curves in the parameter space $(\phi,\tau)$.
IV.3 Asymptotic dynamics time scale
As pointed out in Section IV.1, the Perron-Frobenius analysis provides useful insights into the asymptotic dynamics of the system. Related to our assumption of a finite observation time, this naturally leads to the question: What is the time scale over which such asymptotic behavior emerges? Such a time scale is governed by the spectral properties of the Perron-Frobenius operator (IV.1), in particular by the spectral gap $\Delta\equiv 1-|\mu_{\rm PF}|$, where we recall $\mu_{\rm PF}$ is the largest-modulus eigenvalue, according to which the asymptotic dynamics time scale can be estimated as $t_{\rm as}\sim\Delta^{-1}$.
@@ IV.3 Asymptotic dynamics time scale
Because the Perron-Frobenius operator depends on the measurement period $\tau$, the asymptotic time scale $t_{\rm as}$ varies with $\tau$ as well, as shown in Fig. 5(a). In the limit of small $\tau$, $t_{\rm as}$ diverges due to the quantum Zeno effect, which prevents the walker from reaching—and thus being detected at—the measurement site. As $\tau$ increases, $t_{\rm as}$ monotonically decreases, reaching its minimum at $\tau_{\rm opt}$ (minimum of $|\mu_{\rm PF}|$), after which it quickly increases. Upon crossing the threshold $\tau^{*}$, the system transitions to a regime characterized by a high sensitivity to small variations in the model parameters $\tau$ and $\phi$ [Fig. 2], a sensitivity also reflected in the values of $t_{\rm as}$ [Fig. 5(a)]. This behavior is consistent upon varying the number of sites $N$, with larger $N$ leading to higher values of $t_{\rm as}$. As anticipated at the end of Section IV.1, the $\tau_{\rm opt}$ that minimizes $|\mu_{\rm PF}|$ approaches $\pi/2$ as $N$ increases, a trend that is reflected in the corresponding behavior of the minimum of $t_{\rm as}$ in Fig. 5(a).
The asymptotic time at which the detection probability saturates increases as $N$ increases. However, in practical cases, the experimentalist may be satisfied with the detection probability exceeding a certain threshold. This motivates the study in Fig. 5(b), where $P_{\rm det}$ is shown as a function of $N$ and the total observation time $T$ at $\tau_{\rm opt}$ and $\phi_{\rm opt}$ derived from the Perron-Frobenius analysis. First, we observe that the larger $N$, the larger $T$ at which $P_{\rm det}$ saturates (i.e., $T\sim t_{\rm as}$), and the larger the minimum $T$ at which $P_{\rm det}$ is non-vanishing. Second, we note that the observation time $T$ required to reach, e.g., $P_{\rm det}\approx 80\%$ is almost an order of magnitude lower than that required to saturate. Third, we note that the observation time $T=200$ considered in the present work is sufficiently long to observe the asymptotic dynamics and the saturation of $P_{\rm det}$ for $N<21$, with $T\sim t_{\rm as}$ at $N=21$.
For larger $N$, the system takes longer to enter the asymptotic regime. Provided that the observation time satisfies $T\gtrsim t_{\rm as}$, then $P_{\rm det}$ can saturate and the optimal sampling period $\tau_{\rm opt}$ is identified by the minimum of $|\mu_{\rm PF}|$. Instead, when $T<t_{\rm as}$, the optimal sampling period $\tau_{\rm opt}$ differs from that determined from the Perron-Frobenius analysis (minimum of $|\mu_{\rm PF}|$), say $\tau_{\rm opt}^{\rm(PF)}$, and it decreases with $N$ at a rate that depends on $T$: the longer $T$, the slower the decrease, with $\tau_{\rm opt}$ approaching $\tau_{\rm opt}^{\rm(PF)}$ as $T\to\infty$ (data not shown).
@@
Figure 5: (a) Asymptotic time $t_{\rm as}$ as a function of the sampling period $\tau$ for various odd $N$. (b) Density plot of $P_{\rm det}$ as a function of odd $N$ and total observation time $T$, evaluated at the $\tau_{\rm opt}$ corresponding to each $N$ and at $\delta=(N-1)/2$. In both panels the phase is fixed at $\phi_{\rm opt}=\pi/2N$.
V Conclusions
In this work, we have investigated the beneficial interplay between chirality and local monitoring in enhancing excitation transfer on a ring, within the framework of continuous-time quantum walks. Our findings address the gap in understanding how chirality, $\phi$, and sampling period, $\tau$, can be jointly optimized to overcome the limitations of purely unitary dynamics, usually stemming from destructive interference and presence of dark states. The resulting transfer protocol is optimal, significantly improving transfer probability, and robust, tolerating imperfect parameter tuning. This represents a substantial improvement over methods relying on precise fine-tuning, e.g., the evolution time in continuous-time quantum search [50], a common experimental challenge. Our approach in optimizing the transfer protocol leverages (i) a careful analysis of dark states to identify those, if any, relevant to the dynamics under investigation, and (ii) the Perron-Frobenius spectrum of the non-unitary dynamics to identify the optimal parameters, thus providing a general method for improving transport in monitored quantum systems.
@@ V Conclusions
Beyond deepening the fundamental understanding of first-detected-passage time problems and transport in closed loops, our results offer practical guidelines for designing quantum protocols that require robust and efficient state transfer [51], such as quantum routers [29]. Our study holds the potential to contribute to better design for quantum-walk-based computing [52] and for robust quantum search strategies [53].
Finally, hitting times are expected to directly impact the efficiency of quantum technologies, and the recent implementation of local monitoring on quantum hardware via mid-circuit measurements—either strong [54, 41] or weak [55]—highlights the experimental feasibility of such schemes. Furthermore, restart strategies are emerging as a further option to speed up quantum hitting times in monitored quantum systems [56, 57, 42, 58], offering potential insights for the understanding and design of efficient quantum algorithms.
Acknowledgements.
@@
S.F., L.R., and G.B. acknowledge support by INFN through the project “QUANTUM”. S.F. and G.B. acknowledge support by Grant No. PNRR D. M. 118/2023. This work has been done under the auspices of GNFM-INdAM and M.G.A.P. acknowledges partial support by MUR and EU through the projects PRIN22-PNRR-P202222WBL-QWEST, and NQSTI-Spoke2-BaC QMORE (contract n. PE00000023-QMORE).
Figure 6: Instantaneous transfer probability $P_{\delta}$ from site 00 to $\delta$ as a function of the phase $\phi$ and time $t$ (a) for $N=20$ and $\delta=N/2$ and (b) for $N=21$ and $\delta=(N-1)/2$.
Appendix A Purely-coherent transport under unitary evolution
As anticipated in the main text, the purely-coherent transport under unitary dynamics is not efficient. The instantaneous probability $P_{\delta}(\phi,t)=|\langle\delta|\hat{U}(\phi,t)|\psi_{0}\rangle|^{2}$ typically remains low throughout the evolution, punctuated by narrow, sharp peaks. This is evident from the comparison between Figures  2 and 6: Fig. 6 shows $P_{\delta}$ as a function of the phase $\phi$ and time $t$, revealing that its values are consistently and significantly lower over the entire parameter range considered than those of the detection probability $P_{\rm det}$ in Fig. 2.
Appendix B Perron-Frobenius analysis for even $N$
The Perron-Frobenius analysis in Section  IV.1 focused on the case of odd $N$. Here we discuss the case of even $N$, which needs to be considered more carefully. The main differences between even and odd $N$ root back to the degeneracies of the energy spectrum in Eq. (3). For odd $N$, all the energy levels but one—the highest ($j=0$)—have a two-fold degeneracy when $\phi=0$. These degeneracies are lifted when $\phi\neq 0$. For even $N$, instead, all the energy levels but two—the lowest ($j=N/2$) and the highest ($j=0$)—have a two-fold degeneracy when $\phi=0$. Nonzero phases lift these degeneracies except when $\phi=\pm\pi/N$, in which case all the energy levels, including the lowest and the highest, have a two-fold degeneracy. Following [38], we can construct a dark state from each degenerate level $\lambda_{j}$ as follows
$|\gamma_{j}\rangle=\sqrt{\frac{N}{2}}\Big{(}\langle\delta|\lambda_{j,2}\rangle|\lambda_{j,1}\rangle-\langle\delta|\lambda_{j,1}\rangle|\lambda_{j,2}\rangle\Big{)}.$ (14)
For $\phi=0$, the dark states read [38]
$|\gamma_{j}\rangle=\sqrt{\frac{2}{N}}\sum_{k=0}^{N-1}\sin\left(\frac{2\pi jk}{N}\right)|k\rangle$ (15)
with $j=1,\ldots,N/2-1$, and have no overlap with the initial state, $\langle 0|\gamma_{j}\rangle=0\,\forall j$. As a result, these states do not contribute to the walker’s dynamics and can be considered irrelevant in the evolution. These dark states are eigenstates of the Perron-Frobenius operator with unit-modulus eigenvalue [see Fig. 7(a,b)]. Consequently, the asymptotic dynamics relevant to excitation transfer will be governed by the subleading eigenvalue, defined as $\tilde{\mu}_{\rm PF}\equiv\max_{|\mu_{k}|<1}(\mu_{k})$, i.e., the largest eigenvalue in modulus strictly less than one.
For $\phi=-\pi/N$, instead, the dark states read
$|\gamma_{j}\rangle=\frac{1}{\sqrt{2N}}\sum_{k=0}^{N-1}\left(e^{i\frac{2\pi jk}{N}}+e^{-i\frac{2\pi(j+1)k}{N}}\right)|k\rangle$ (16)
with $j=0,\ldots,N/2-1$, and always overlap with the initial state, $\langle 0|\gamma_{j}\rangle=\sqrt{2/N}\,\forall j$. The initial state turns out to be completely dark for $\phi=-\pi/N$, as it is the equal superposition of all the $N/2$ dark states forming a basis for the dark subspace, $|0\rangle=\sqrt{2/N}\sum_{j=0}^{\frac{N}{2}-1}|\gamma_{j}\rangle$. The case $\phi=\pi/N$ can be treated analogously. Therefore, $P_{\rm det}(n\to\infty)=1$ if $\phi=0$, and $P_{\rm det}(n\to\infty)=0$ if $\phi=\pm\pi/N$ [23].
Figure 7: Eigenvalues of the Perron-Frobenius operator (IV.1) on the unit circle for two representative values of the sampling period, (a) $\tau<\tau^{*}$ and (b) $\tau>\tau^{*}$ at $\phi=0$. (c) Density plot of the largest-modulus PF eigenvalue $|\mu_{\rm PF}|$ as a function of $\phi$ and $\tau$. (d) Curve $|\mu_{\rm PF}|$ as a function of $\tau$ at fixed $\phi=\pi/2N$. (e) Curve $|\mu_{\rm PF}|$ as a function of $\phi$ at fixed $\tau=1.53$. The red circle denotes the subleading PF eigenvalue $|\tilde{\mu}_{\rm PF}|$, which is the relevant one for the asymptotic dynamics at $\phi=0$ (the initial state has zero overlap with dark states). Other parameters are: $N=20$, $\delta=10$.
The study of the largest-modulus Perron-Frobenius eigenvalue $\mu_{\rm PF}$ is shown in Fig. 7(c). Similarly to the case of odd $N$, $|\mu_{\rm PF}|$ smoothly attains a local minimum (which we identify as $\tau_{\rm opt}$) in the interval $\tau<\tau^{\star}$, and then rapidly oscillates [Fig. 7(d)]. As for the dependence on the phase $\phi$, $\mu_{\rm PF}$ shows two minima at $\phi\neq 0$, in contrast with the optimal phase $\phi_{\rm opt}=0$ we expect for even $N$ from the numerical optimization reported in the main text (Section III). On the other hand, as discussed above, dark states built using degenerate energy levels are irrelevant for excitation transfer at $\phi=0$. At $\phi=0$, one should thus consider the modulus of the subleading eigenvalue, $|\tilde{\mu}_{\rm PF}|$ [red dot in Fig. 7(e)], which is lower than the two local minima of $|\mu_{\rm PF}|$ when $\phi\neq 0$. We therefore conclude that the optimal phase for even $N$ is $\phi=0$.
Appendix C Density of dark states increases with $\tau$
In this Appendix, we prove that the density of dark states—i.e., the number of dark states in a certain interval $[\phi,\phi+\Delta\phi]$ at fixed $\tau$—grows with $\tau$ in the region $\tau>\pi/2$. As discussed in Sec IV.2, outside of the degenerate case, dark states correspond to solutions of Eq. (13). Graphically, these solutions are the intersections of two functions, $y_{1}=k\pi$ and $y_{2}=2\tau\sin\left[\pi(m-n)/N\right]\sin\left[\phi-\pi(m+n)/N\right]$. On the plane $(\phi,y)$, the function $y_{1}$ represents horizontal lines at integer multiples of $\pi$; at fixed $\tau,m,n,N$, the function $y_{2}$ is a sinusoidal function in $\phi$, with $\tau$ acting as an amplification factor, provided that $m,n$ satisfy $\sin\left[\pi(m-n)/N\right]\neq 0$. It is straightforward to see that the number of horizontal lines $y_{1}=k\pi$ crossed by the function $y_{2}$ grows in any interval $[\phi,\phi+\Delta\phi]$ as $\tau$ increases, as was to be demonstrated. In other words, for larger values of $\tau$, a finer tuning on $\phi$ is needed to avoid dark states, solutions of $y_{1}=y_{2}$. Searching for a robust transfer protocol, this further motivates our choice of restricting to $\tau<\tau^{*}$.

### v2

@@ Abstract
$\hat{U}(\tau)|\gamma_{n,m}\rangle=\sqrt{\frac{N}{2}}\Big{(}e^{-i\lambda_{m}\tau}\langle\delta|\lambda_{n}\rangle$ $|\lambda_{m}\rangle$
$-e^{-i\lambda_{n}\tau}\langle\delta|\lambda_{m}\rangle$ $|\lambda_{n}\rangle\Big{)},$ (8)
@@
in which case the system never visits |$\ket{\delta}$, resulting in $P_{\rm det}(\infty)=0$. If no pairs of eigenvalues $\lambda_{m},\lambda_{n}$ satisfy the condition (9) and $\langle\lambda_{j}|\delta\rangle\neq 0$ $\forall j$, then the Hilbert space is fully bright, and there are no dark states [38]. Note that when Eq. (9) holds true, we have that $[\mathbb{I}-\hat{D}]\hat{U}(\tau)|\gamma_{n,m}\rangle\,=\,e^{-i\lambda_{m}\tau}|\gamma_{n,m}\rangle$, i.e., these dark states are eigenstates of the Perron-Frobenius operator $\hat{O}$ and are the only ones with unit-modulus eigenvalue  [^4] The time-evolution operator $\hat{U}(\tau)$ in Eq. (3) is unitary, thus preserves the norm and has unit-modulus eigenvalues. Accordingly, for the Perron-Frobenius operator (Optimal quantum transport on a ring via locally monitored chiral quantum walks) to have an eigenvalue of modulus 1 with corresponding eigenstate $|\mu\rangle$, it must be that $\hat{D}\hat{U}(\tau)|\mu\rangle=0$; otherwise, the non-unitary operator $(\mathbb{I}-\hat{D})$ would reduce the norm of $\hat{U}(\tau)|\mu\rangle$, preventing a unit-modulus eigenvalue of $\hat{O}$. In other words, $|\mu\rangle$ is eigenstate of $\hat{O}$ with unit-modulus eigenvalue only if $\hat{O}|\mu\rangle=\hat{U}(\tau)|\mu\rangle$, meaning that $|\mu\rangle$ must be an eigenstate of $\hat{U}(\tau)$ orthogonal to $|\delta\rangle$.. In our model the condition $\langle\lambda_{j}|\delta\rangle\neq 0\;\forall j$ is satisfied by any localized state $|\delta\rangle$ [see Eq. (2)], as our detection states, while the condition (9) can be satisfied in two ways:
@@
For even $N$, when $\phi=0$ there exist $\frac{N}{2}-1$ two-fold degenerate levels. In this case, all the stationary dark states constructed from the pairs of eigenstates in each degenerate level are orthogonal to the initial state. Therefore, they neither contribute to the dynamics of the system nor affect the detection probability. When $\phi=\pm\pi/N$ there exist $N/2$ two-fold degenerate levels. It can be proved (see Sec. S2 of SM [44]) that the initial state lies in the dark subspace spanned by the stationary dark states constructed from these degenerate levels. As a consequence, in this case the walker is never detected in $|\delta\rangle$, yielding $P_{\rm det}(\pm\pi/N)=0$ in Fig. 2(a).
For odd $N$, calculations analogous to those in Sec. S2 of SM [44] reveal that for $\phi=0,\pm\pi/N$ the total overlap of the initial state with the dark space is $1/2$, thus explaining the minima of the detection probability for these values of $\phi$ in Fig. 2(b). For such values of $\phi$, there exist $(N-1)/2$ two-fold degenerate levels, each contributing a dark state that, together, form the basis spanning the dark space.
(ii) Using any nondegenerate couple of eigenvalues $\lambda_{m}\neq\lambda_{n}$, there may exist proper combinations of $\tau$ and $\phi$ that lead to $\lambda_{m}\tau=\lambda_{n}\tau+2k\pi$ for some $k\in\mathbb{Z}$. Plugging the eigenvalues (2) in the latter expression, i.e., in Eq. (9), we derive the condition relating $\tau$ and $\phi$ under which dark states appear
@@
Such a lower bound is exactly attained for even $N$ when $\phi=0$. In all other cases, the $\tau$ at which the first dark state appears approaches $\pi/2$ as $N$ grows. This is because the discrete arguments of the sine functions in Eq. (10) become denser in the interval $[0,2\pi)$ for increasing $N$. For large $N$, it is therefore possible to find pairs $(m,n)$ such that the product of the two sine functions, regardless of $\phi$, is arbitrarily close to 1, implying the condition $\tau\gtrsim\pi/2$. In other words, for any $\phi$ the first dark state associated with nondegenerate levels will arise at $\tau\to\pi/2$ as $N\to\infty$, appearing as a dark-state line at $\tau\to\pi/2$ in the parameter space $(\phi,\tau)$. In general, for finite $N$, solutions to Eq. (10) constitute dark-state curves in the $(\phi,\tau)$-space, which correspond to low-detection-probability curves. Figures 2(a,b) represent a clear example: The detection probability in the parameter space $(\phi,\tau)$ is low along the curves associated with the presence of dark states, including a nearly straight line at $\tau\gtrsim\pi/2$. In the region $\tau\geq\pi/2$, the dark-state curves become denser as $N$ increases, because the number of dark states from nondegenerate levels is of order $O(N^{2})$, and also as $\tau$ increases (see Sec. S3 of SM [44]).
@@
To summarize, in the region $\tau<\pi/2$, for even $N$ dark states at $\phi=0$ do not affect the evolution of $|\psi_{0}\rangle$ yielding $P_{\rm det}(\infty)=1$, while for $\phi=\pm\pi/N$ the initial state is a dark state, yielding $P_{\rm det}(\infty)=0$; for odd $N$, the initial state has a finite overlap with dark states at $\phi=0,\pm\pi/N$, yielding $P_{\rm det}(\infty)=1/2$. For $\phi\neq 0,\pm\pi/N$, the Hilbert space is fully bright in this region, so $P_{\rm det}(\infty)=1$ for both even and odd $N$. On the other hand, in the region $\tau\geq\pi/2$, the number of dark states increases as $N$ and $\tau$ increase, yielding low-detection-probability curves in the $(\phi,\tau)$-space.
@@
Time scale of the asymptotic dynamics.—As pointed out above, the Perron-Frobenius analysis provides useful insights into the asymptotic dynamics of the system. Related to our assumption of a finite observation time, this naturally leads to the question: What is the time scale over which such asymptotic behavior emerges? Such a time scale is governed by the spectral properties of the Perron-Frobenius operator (Optimal quantum transport on a ring via locally monitored chiral quantum walks), in particular by the spectral gap $\Delta\equiv 1-|\mu_{\rm PF}|$, where we recall $\mu_{\rm PF}$ is the largest-modulus eigenvalue, according to which the asymptotic dynamics time scale can be estimated as $t_{\rm as}\sim\Delta^{-1}$. Because the Perron-Frobenius operator depends on the detection period $\tau$, the asymptotic time scale $t_{\rm as}$ varies with $\tau$ as well, as shown in Fig. 5(a). In the limit of small $\tau$, $t_{\rm as}$ diverges due to the quantum Zeno effect, which prevents the walker from reaching—and thus being detected at—the measurement site. As $\tau$ increases, $t_{\rm as}$ monotonically decreases, reaching its minimum at $\tau_{\rm opt}$ (minimum of $|\mu_{\rm PF}|$), after which it quickly increases. Upon crossing the threshold $\tau^{*}$, the system transitions to a regime characterized by a high sensitivity to small variations in the model parameters $\tau$ and $\phi$ [Fig. 2], a sensitivity also reflected in the values of $t_{\rm as}$ [Fig. 5(a)]. This behavior is consistent upon varying the number of sites $N$, with larger $N$ leading to higher values of $t_{\rm as}$. The fact that the $\tau_{\rm opt}$ minimizing $|\mu_{\rm PF}|$ approaches $\pi/2$ as $N$ increases is reflected in the corresponding behavior of the minimum of $t_{\rm as}$ in Fig. 5(a).
@@
Figure 5: (a) Asymptotic time $t_{\rm as}$ as a function of the detection period $\tau$ for various odd $N$. (b) Density plot of $P_{\rm det}$ as a function of odd $N$ and total observation time $T$, evaluated at the $\tau_{\rm opt}$ corresponding to each $N$ and at $\delta=(N-1)/2$. In both panels the phase is fixed at $\phi_{\rm opt}=\pi/2N$.
The asymptotic time at which the detection probability saturates increases as $N$ increases. However, in practical cases, the experimentalist may be satisfied with the detection probability exceeding a certain threshold. This motivates the study in Fig. 5(b), where $P_{\rm det}$ is shown as a function of $N$ and the total observation time $T$ at $\tau_{\rm opt}$ and $\phi_{\rm opt}$ derived from the Perron-Frobenius analysis. First, we observe that the larger $N$, the larger $T$ at which $P_{\rm det}$ saturates (i.e., $T\sim t_{\rm as}$), and the larger the minimum $T$ at which $P_{\rm det}$ is non-vanishing. Second, we note that the observation time $T$ required to reach, e.g., $P_{\rm det}\approx 80\%$ is almost an order of magnitude lower than that required to saturate. Third, we note that the observation time $T=200$ considered in the present work is sufficiently long to observe the asymptotic dynamics and the saturation of $P_{\rm det}$ for $N<21$, with $T\sim t_{\rm as}$ at $N=21$. For larger $N$, the system takes longer to enter the asymptotic regime. Provided that the observation time satisfies $T\gtrsim t_{\rm as}$, then $P_{\rm det}$ can saturate and the optimal detection period $\tau_{\rm opt}$ is identified by the minimum of $|\mu_{\rm PF}|$. Instead, when $T<t_{\rm as}$, the optimal detection period $\tau_{\rm opt}$ differs from that determined from the Perron-Frobenius analysis (minimum of $|\mu_{\rm PF}|$), say $\tau_{\rm opt}^{\rm(PF)}$, and it decreases with $N$ at a rate that depends on $T$: the longer $T$, the slower the decrease, with $\tau_{\rm opt}$ approaching $\tau_{\rm opt}^{\rm(PF)}$ as $T\to\infty$ (data not shown).
Conclusions.—In this Letter, we have investigated how chirality and local monitoring jointly enhance excitation transfer, modeled as a continuous-time quantum walk on a ring. We show that optimizing both the chiral phase $\phi$ and the detection period $\tau$ overcomes limitations of purely unitary dynamics, such as destructive interference and dark states. The resulting transfer protocol is both optimal, yielding significantly higher detection probabilities, and robust, requiring no fine-tuning of parameters or tailored initial-state preparations. Our approach combines two key insights: (i) The identification of dynamically relevant dark states (see Sec. S2 of SM [44]), and (ii) the spectral analysis of the non-unitary Perron-Frobenius operator to determine optimal parameters. This provides a general framework for improving transport in monitored quantum systems.
@@
Beyond advancing the understanding of first-detected-passage problems in closed loops, our results offer practical guidelines that can be readily extended beyond quantum transport to the design of efficient quantum protocols in domains where destructive interference and dark states usually limit performance, such as search algorithms [55, 56], state transfer [57], quantum routers [29], and potentially in quantum-walk-based computing [58].
Finally, hitting times play a central role in the performance of quantum protocols, as they directly affect how rapidly a target state can be reached or detected. The recent implementation of local monitoring via mid-circuit measurements—either strong [59, 41] or weak [60]—demonstrates the growing experimental feasibility of accessing quantum hitting-time statistics on current quantum hardware. In parallel, restart strategies have emerged as a new, powerful control mechanism to further speed up hitting times in monitored quantum systems [61, 62, 42, 63]. Together, these advances underscore the relevance of our approach for designing efficient and experimentally viable quantum protocols.
@@
Acknowledgments.—S.F., G.B., and L.R. acknowledge support by INFN through the project “QUANTUM”. S.F. and G.B. acknowledge support by Grant No. PNRR D. M. 118/2023. This work has been done under the auspices of GNFM-INdAM and M.G.A.P. acknowledges partial support by MUR and EU through the projects PRIN22-PNRR-P202222WBL-QWEST, and NQSTI-Spoke2-BaC QMORE (contract n. PE00000023-QMORE).